OpenStudy (anonymous):

Write the equation of the line that passes through (2, 5) with a slope of 3 in point-slope form. y - 2 = 3(x - 5) y + 8 = 3(x + 5) y - 2 = 3(x - 1) y - 5 = 3(x - 2)

OpenStudy (whpalmer4):

Point-slope form is \[y-y_0 = m(x-x_0)\] for a line with slope \(m\) passing through point \((x_0,y_0)\)

OpenStudy (anonymous):

y-5= 3 (x-2)

OpenStudy (anonymous):

like this.. y-y1 = m (slope) (x-x1)